OpenStudy (anonymous):

what is 27 x 1/3? help please

OpenStudy (timbo5lice):

9

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you should do : \[\frac{ 27 }{ 1 }*\frac{ 1 }{ 3 }=\frac{ 27 }{ 3 }=9\]

OpenStudy (anonymous):

thanks im closing now.